1.Determine the nature of the stationary value of x= t^3 - 3t + ty^2

2. A rectangle box whose value is 32 is open at the top,if the surface area is 2(L+B)H + LB where LBH are Length,Breadth and Height Respectively

i.Find the Dimension of the box that may require list material

ii.investigate weather the dimension found requires List material.


1
Expert's answer
2021-12-06T16:00:28-0500

1.


"x(t, y)=t^3 - 3t + ty^2"


"\dfrac{\partial x}{\partial t}=3t^2-3+y^2"

"\dfrac{\partial x}{\partial y}=2yt"

Find the critical point(s)


"\dfrac{\partial x}{\partial t}=0,"

"\dfrac{\partial x}{\partial y}=0"


"3t^2-3+y^2=0"

"2yt=0"

"t=0, y=-\sqrt{3}"

"t=0, y=\sqrt{3}"

"y=0,t=-1"

"y=0, t=1"

"\dfrac{\partial^2 x}{\partial t^2}=6t"

"\dfrac{\partial^2 x}{\partial t\partial y}=2y"

"\dfrac{\partial^2 x}{\partial y^2}=2t"

"D=\begin{vmatrix}
 6t & 2y \\
 2y & 2t
\end{vmatrix}=12t^2-4y^2"

"t=0, y=-\sqrt{3}"


"\dfrac{\partial^2 x}{\partial t^2}|_{t=0, y=-\sqrt{3}}=0"

"D|_{t=0, y=-\sqrt{3}}=-12<0"

"x(0, -\sqrt{3})" is not a local maximum or local minimum.



"t=0, y=\sqrt{3}"

"\dfrac{\partial^2 x}{\partial t^2}|_{t=0, y=\sqrt{3}}=0"

"D|_{t=0, y=\sqrt{3}}=-12<0"

"x(0, \sqrt{3})" is not a local maximum or local minimum.



"t=-1, y=0"

"\dfrac{\partial^2 x}{\partial t^2}|_{t=-1, y=0}=-6<0"

"D|_{t=-1, y=0}=12>0"

"x(-1, 0)" is a local maximum.



"t=1, y=0"

"\dfrac{\partial^2 x}{\partial t^2}|_{t=1, y=0}=6>0"

"D|_{t=1, y=0}=12>0"

"x(1, 0)" is a local minimum.


2.

Given

"LBH=32=>H=\dfrac{32}{LB}"

"S=(2L+2B)H+LB"

Substitute


"S(L, B)=(2L+2B)(\dfrac{32}{LB})+LB"

"=\dfrac{64}{B}+\dfrac{64}{L}+LB, L>0, B>0"


"\dfrac{\partial S}{\partial L}=-\dfrac{64}{L^2}+B"

"\dfrac{\partial S}{\partial B}=-\dfrac{64}{B^2}+L"

Find the critical point(s)


"\dfrac{\partial S}{\partial L}=0,"

"\dfrac{\partial S}{\partial B}=0"


"-\dfrac{64}{L^2}+B=0"

"-\dfrac{64}{B^2}+L=0"

"BL^2=64=L^2B"

"B=L=4"

"\dfrac{\partial^2 S}{\partial L^2}=\dfrac{128}{L^3}"

"\dfrac{\partial^2 S}{\partial B^2}=\dfrac{128}{BL^3}"

"\dfrac{\partial^2 S}{\partial L\partial B}=1"

"D=\begin{vmatrix}
128/L^3 & 1 \\
 1 & 128/B^3
\end{vmatrix}=16384/(LB)^3-1"

"\dfrac{\partial^2 S}{\partial L^2}|_{L=4, B=4}=2>0"

"D|_{L=4, B=4}=4-1=3>0"

"S(4, 4)" is a local minimum.Since the function "S(L,B)" has the only extreme value for "L>0, B>0," then S(4, 4) is the absolute minimm for "L>0, B>0."


"H=\dfrac{32}{4(4)}=2"

The dimensions of the box for least material are

"Length\times Breadth\times Height"

"=4\times4\times 2"